# Growth rate of liquidity provider’s wealth in G3Ms

### v3

@@ 4.3. Time-homogeneous reflected diffusion
This corollary, which can also be found in [GW13], provides a concise expression for the long-term average behavior of the processes $L_{t}$ and $U_{t}$, which regulate the reflection of the mispricing process at the boundaries.
@@
of a constant rebalanced portfolio with the same weight, as discussed in [Fer02, §1.1]. Furthermore, the term $\frac{(1-\gamma)w(1-w)}{1-w+\gamma w}\alpha+\frac{(1-\gamma)w(1-w)}{\gamma(1-w)+w}\beta$ converges to the excess growth rate $\frac{w(1-w)}{2}\sigma^{2}$ as the fee tier $\gamma$ approaches 1.
@@ 4.4. Optimal fees and optimal growth
This result is consistent with the calculation in [Har13, Proposition 6.6], which utilizes the stationary distribution of the mispricing process.
@@ 4.4.1. Numerical Analysis and Optimal Fees
We now numerically investigate the growth rate of LP wealth in G3Ms across different fee tiers $\gamma$ and asset weights $w$. This analysis extends the work in [TW20, §5], which focused on the specific case of $w=\frac{1}{2}$.
@@ Conclusion
This paper investigated the growth of liquidity provider (LP) wealth in Geometric Mean Market Makers (G3Ms), explicitly considering the impacts of continuous-time arbitrage and transaction fees. We extended the analysis of arbitrage in LP profitability beyond the constant product models studied in [TW20] to encompass a broader class of G3Ms.
Our model-free approach, which builds on the framework in [MMRZ22b, MMRZ22a, MMR23], provides a refined understanding of G3M dynamics under the influence of arbitrage and fees. A key result is our characterization of LP wealth growth through a stochastic model driven by the mispricing process between the G3M and an external reference market. This model elucidates the crucial role of trading fees in LP returns and demonstrates how arbitrage activity establishes a no-trading band around the reference market price.
Our analysis reveals that the adverse selection risk posed by arbitrageurs necessitates a nuanced understanding of LP wealth dynamics. Notably, we demonstrated that, under certain market conditions, G3Ms with fees can outperform both buy-and-hold and constant rebalanced portfolio strategies. This finding highlights the potential of G3Ms as a competitive investment product within the DeFi landscape, effectively functioning as on-chain index funds.
Several promising avenues for future research emerge from our work. These include incorporating noise trader order flows to create more realistic market dynamics, investigating the interplay between different AMM liquidity pools, refining LP wealth growth models to account for factors like impermanent loss, and extending the analysis to G3Ms with dynamic weights, drawing connections to stochastic portfolio theory (SPT) as explored in [Eva21, Fer02, KF09]. These research directions will further enrich our understanding of AMM functionality and its broader implications for liquidity providers and the DeFi ecosystem.
@@ Acknowledgement
C. Y. Lee was supported in part by a research startup fund of the Chinese University of Hong Kong, Shenzhen and the Shenzhen Peacock fund 2025TC0013.
@@ Appendix C Time-inhomogeneous reflected diffusion
In this appendix, we provide the proof of the long-term time averaged expectation of a time-inhomogeneous reflected diffusion as stated in Theorem 4.13. For fixed $t,x$, we shall sometimes suppress the reference to $t$, $x$ in the transition density $p$ and simply denote $p(s,y|t,x)$ by $p(s,y)$ for simplicity. For any function $\varphi$ defined in [c,c][-c,c], $\|\varphi\|_{2}$ denotes the $L^{2}$ norm of $\varphi$ in [c,c][-c,c]. We start with stating an estimate of the $L^{2}$-norm between the transition density $p$ and the stationary density $q$ in the following lemma, whose proof is omitted (for interested readers, we refer it to, for instance, [Kah83], see (3.21) on P.276), is classical and crucial to the proof that follows.

### v4

@@ 4.3. Time-homogeneous reflected diffusion
This corollary, which can also be found in [glynn2013central], provides a concise expression for the long-term average behavior of the processes $L_{t}$ and $U_{t}$, which regulate the reflection of the mispricing process at the boundaries.
@@
of a constant rebalanced portfolio with the same weight, as discussed in [fernholz2002stochastic, §1.1]. Furthermore, the term $\frac{(1-\gamma)w(1-w)}{1-w+\gamma w}\alpha+\frac{(1-\gamma)w(1-w)}{\gamma(1-w)+w}\beta$ converges to the excess growth rate $\frac{w(1-w)}{2}\sigma^{2}$ as the fee tier $\gamma$ approaches 1.
@@ 4.4. Optimal fees and optimal growth
This result is consistent with the calculation in [harrison2013brownian, Proposition 6.6], which utilizes the stationary distribution of the mispricing process.
@@ 4.4.1. Numerical Analysis and Optimal Fees
We now numerically investigate the growth rate of LP wealth in G3Ms across different fee tiers $\gamma$ and asset weights $w$. This analysis extends the work in [tassy2020growth, §5], which focused on the specific case of $w=\frac{1}{2}$.
@@ Conclusion
This paper studied the long-term growth of liquidity provider (LP) wealth in Geometric Mean Market Makers (G3Ms), explicitly account for the effects of continuous-time arbitrage and transaction fees. Building on the framework of [milionis2022quantifying, milionis2022automated, milionis2023automated] and extending the constant product analysis of [tassy2020growth] to a broader class of G3Ms, we developed a tractable stochastic model driven by the mispricing process between the G3M pool and an external reference market.
A central finding is that trading fees play a dual role in LP wealth dynamics: they generate income through the no-arbitrage band they impose on the pool price, while simultaneously mediating the adverse selection risk posed by arbitrageurs. Our explicit formulas for the long-term logarithmic growth rate of LP wealth capture both effects, and reveal that — under appropriate fee selection — G3Ms can outperform both buy-and-hold and constant-rebalanced portfolio strategies. The result connects G3M dynamics to the excess growth rate in the Stochastic Portfolio Theory [fernholz2002stochastic], and supports the view of G3Ms as viable on-chain index fund infrastructure: short-term arbitrage losses are transformed, through the fee mechanism, into long-run volatility harvesting.
The present analysis is carried out in the two-asset setting, which allows for a complete and tractable treatment of the mispricing dynamics and their ergodic properties. The structural form of these dynamics, however, extends naturally to the multi-asset case. For a pool with $n$ assets, the log-mispricing of asset $i$ relative to asset $j$ satisfies
$dZ_{t}^{ij}=d\ln S_{t}^{ij}+\sum_{k\neq i}\alpha^{ij}_{ik}dL_{t}^{ik}-\sum_{k\neq j}\beta^{ij}_{kj}dU_{t}^{kj},$
where $\alpha^{ij}_{ik}$ and $\beta^{ij}_{kj}$ are constants determined by the fee parameter $\gamma$, the pool weights $w$, and the normalization of order flows, and $L_{t}^{ij}$, $U_{t}^{ij}$ are continuous non-decreasing processes that increase only when $Z_{t}^{ij}=\ln\gamma$ and $Z_{t}^{ij}=-\ln\gamma$, respectively. The principal challenge in extending our results to this setting lies not in the dynamics’ structural form, which carries over directly, but in characterizing the ergodic behavior of the boundary local times $L_{t}^{ij}$ and $U_{t}^{ij}$. These are driven by a higher-dimensional reflected diffusion for which the existing literature does not yet provide a complete theory. We intend to pursue this direction in subsequent work.
Several further directions emerge naturally from this work. These include incorporating noise-trader order flow to capture the contribution of uninformed trading to LP profitability, investigating the interplay between multiple AMM liquidity pools and a fragmented market microstructure, and extending the framework to G3Ms with dynamic weights in the spirit of [evans2021liquidity, fernholz2002stochastic, karatzas2009stochastic]. Developing the ergodic theory for multi-asset reflected diffusions and deriving the corresponding LP wealth growth rates for pools with $n>2$ assets constitutes the most immediate and natural direction for future research, which we intend to pursue in subsequent work.
@@ Acknowledgement
C. Y. Lee was supported in part by a research startup fund of the Chinese University of Hong Kong, Shenzhen, and the Shenzhen Peacock fund 2025TC0013.
@@ Appendix C Time-inhomogeneous reflected diffusion
In this appendix, we provide the proof of the long-term time averaged expectation of a time-inhomogeneous reflected diffusion as stated in Theorem 4.13. For fixed $t,x$, we shall sometimes suppress the reference to $t$, $x$ in the transition density $p$ and simply denote $p(s,y|t,x)$ by $p(s,y)$ for simplicity. For any function $\varphi$ defined in [c,c][-c,c], $\|\varphi\|_{2}$ denotes the $L^{2}$ norm of $\varphi$ in [c,c][-c,c]. We start with stating an estimate of the $L^{2}$-norm between the transition density $p$ and the stationary density $q$ in the following lemma, whose proof is omitted (for interested readers, we refer it to, for instance, [kahane1983asymptotic], see (3.21) on P.276), is classical and crucial to the proof that follows.